A ball is thrown with an initial horizontal velocity of 30 m/s from a height of 3m above the ground and 40m from a vertical wall. How far from the wall will the ball rebound for the first time? 


1
Expert's answer
2021-06-21T05:52:16-0400

"v_x = 30m/s\\
h = 3m\\
x= 40m\\"


"v² = u²+ 2as\\
v_x² = u_x² + 2a_xx\\
v_² = 30² + 0\\
v_x = 30m/s"


"t = \dfrac{x}{v_y}= \dfrac{40}{30} =1.33s"


"y = u_yt+ \frac{1}{2} a_yt²\\
y = 0(1.33) + \frac{1}{2}(10)(1.33)²\\
y = 8.89m"

The ball would have fallen before it got to the wall.